The tile work was done yesterday after noon.
This morning Mike Gerfen and me lowered the enclosure frame to normal height.
Keven- janitor and I removed plastic covers from chambers, racks, SP, MC2 and clean tool boxes.
The afternoon Jenne, Kiwamu, Joe and Aiden cleaned the enclosure inside out. The particle count measured zero inside the enclosure with HEPAs on when the covers were
removed. The MOPA and all other components were happy to see us in excellent condition.
This table height is very user friendly!
Safety grounds were reconnected.
Atm1, new tiles around the concrete slab
Atm2, frame lowered with low cross bars reinstalled
Atm3, the enclosure frame's north west foot is connected to ground
Atm4, PSL optical table is connected to ground at the north east corner through 1 Mohm
Atm5, PSL optical table level at a stimulating, back-friendly height